A car of mass 9500kg is moving up a cliff from rest with a velocity of v= (3t²+2t)i +5tj m/s for 1 minute, if the slope of the cliff is 1, what distance was covered by the car in 1 minute.


1
Expert's answer
2021-01-29T16:57:01-0500
"x(t)=3\frac{t^3}{3}+2\frac{t^2}{2}=t^3+t^2\\y(t)=5t"

"x(60)=60^3+60^2=219600\ m\\y(60)=5(60)=300\ m\\s=\sqrt{219600^2+300^2}=219600\ m"